Sandy held a garage sale during which she charged a dime for everything, but accepted a nickel if the buyer bargained well. At the end of the day she realized she had sold all 12 items and raked in a grand total of 95 cents. She had only dimes and nickels. How many of each did she have? (please show work) Thank you so much: Sandy held a garage sale during which she charged a dime for everything, but accepted a nickel if the buyer bargained well. At the end of the day she realized she had sold all 12 items and raked in a grand total of 95 cents. She had only dimes and nickels. How many of each did she have? (please show work) Thank you so much
Answer by stanbon(18984)

Sandy held a garage sale during which she charged a dime for everything, but accepted a nickel if the buyer bargained well. At the end of the day she realized she had sold all 12 items and raked in a grand total of 95 cents. She had only dimes and nickels. How many of each did she have? (please show work) Thank you so much
-------------------------------------------------
Quantity: d + n = 12
Value: 10d + 5n = 95
------------------------
Multiply the 1st equation by 10 to get:
10d + 10n = 120
10d + 5n = 95
---------------
Subtract 2nd from 1st to get:
5n = 25
n = 5 (# of nickels)
Since d+n + 12, d= 12-5; d = 7 (# of dimes)
===============================================
Cheers,
Stan H.